Presented below are the financial statements of Riverbed Corp. Riverbed Corp Comparative Balance Sheets December 31 Assets 2017 84,000 48,000 48,000 33,600 67,200 48,000 144,000 187,200 (76,800) (57,600) 2016 Cash Accounts receivable Inventory Property, plant, and equipment Accumulated depreciation Total $266,400 259,200 Liabilities and StockholdersEquity Accounts payable Income taxes payable Bonds payable Common stock Retained earnings Total $45,600 36,000 16,80019,200 40,800 79.200 43,200 33,600 91,200 $266,400 259,200 120,000
Income Statement For the Year Ended December 31,2017 Cost of goods sold Gross proft 420,000 160,800 Selling expenses $ 43,200 Administrative expenses 14400 57.600 Income from operations Interest expense Income before Income taxes Income tax expense Net income 7,200 96,000 19,200 $76800 Additional data: 1. Depreciation expense was $ 42,000 2. Dividends declared and paid were $ 48,000. was sold for $ 20,400 cash. This equipment cost $ 43.200 originlly and had accumulated depreciation of s 22.800 at the time of sale
Further analysis reveals the following 1. Accounts payable pertain to merchandise suppliers 2. All operating expenses except for depreclation were paid In cash 3. All depreciation expense is in the selling expense category 4. All sales and purchases are on account. Prepare a statement of cash flows for Riverbed Corp using the direct method.(Show amounts in the investing and financing sections that decrease cash flow with either a-sign e-g.-15,000 or in parenthesis eg.(15,000)) Riverbed Corp Statement of Cash Flows 2 8
media%2Fd2c%2Fd2c717eb-9fee-430b-98b0-47
Compute free cash flow. (Show a negative free cash flow with either a-sign es.-15,000 or in parenthesis eg. (15,000) Free cash flow$ Attempt
Answer #1

Answer:

(A).Preparation of Cash flow statement of Riverbed Corp under direct method :

Cash flow from operating activities = $ 92,400

Cash flow from investing activities = $ 20,400

Cash flow from financing activities = $ (76,800)

Net increase in cash = $ 36,000

Add Opening Cash balance = $ 48,000

Closing cash balance as on Dec 31,2017 = $ 84,000

1 Riverbed Corp Statement of Cash Flows For the year ended December 31,2017 (Direct Method) 4 Particulars Amount Q)Amount ) a

Working Notes:

1.Cash receipt from customers =

Sales + Opening Accounts receivable - Closing Accounts receivable

= 580,800 +33,600 - 48,000 = $566,400

2. Cash payment to suppliers =

Purchases + Opening accounts payable - Closing accounts payable

= 439,200 + 36,000 - 45,600 = $ 429,600

[Purchase = Cost of goods sold + Closing inventory - opening inventory = 420,000 + 67,200 - 48,000 = $439,200]

3. Operating expenses =

Selling expenses + Administrative expenses - Depreciation expense

= 43,200 + 14,400 - 42,000 = $ 15,600

4. Cash payment for income tax expenses =

Income tax expense + opening income tax payable - Closing income tax payable

= 19,200 + 19,200 -16,800 = $ 21,600

5. Redemption of bonds = 79,200 - 40,800 = $38,400

6. Issuance of common stock = 43,200 - 33,600 = $9,600

(B). Calculation of free cash flow

Free Cash Flow =

Net cash flow from operating activities - Capital expenditure - Dividend paid in cash

= $ 92,400 - 0 - 48,000 = $ 44,400